OpenStudy (anonymous):

Evaluate and simplify if possible: for m = 2, n = -3, and p = -1 -6m + 2n - p

OpenStudy (anonymous):

@babygirl121314

OpenStudy (anonymous):

@xiphoidx

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

-6(2)+2(-3)-(-1)=-12-6+1=-18+1=-17

OpenStudy (anonymous):

\[2\times 4(3\times-1-4\times-3) \over 3(5\times4+4\times-3)\]